OpenStudy (anonymous):

Can someone help me through this? A certain region is losing about 15 million acres of rain forest each year. a) find the linear function f that calculates the change in acres of rain forest in millions in x years. b) find f(3) and interpret the result. So.. a) f(x) = b) f(3) =

OpenStudy (pfenn1):

The first year, the change in acres of rain forest is how much?

OpenStudy (anonymous):

15 million

OpenStudy (pfenn1):

Right, and the next year, the change in acres of rain forest is how much?

OpenStudy (anonymous):

the same?

OpenStudy (pfenn1):

Correct. So you can write a linear equation in the form f(x)=kx where you have to figure out what k is. So for year 1 (x=1), f(1)=15 million. For year 2 (x=2), f(2) = 15 million (for the first year) and 15 million (for the second year) or 30 million. so set your equation f(2)=30 million = 2k. What must k equal?

OpenStudy (anonymous):

AreaRegion(t)=(RegionArea(t=0))-(15acres/year)(t) [ this a linear equation A(t)=A0-15*t]

OpenStudy (anonymous):

is f(3)=45million?

OpenStudy (pfenn1):

Correct. What did you get for f(x)=?

OpenStudy (anonymous):

I dont know yet :( F3 was easier, ..LOL .. i am 80% understasnding f(x) ...let me look at it again.

OpenStudy (pfenn1):

well, the pattern is: f(1)=15 million = 1k f(2)=30 million = 2k f(3)=45 million = 3k etc So what must k equal for these equations to be true?

OpenStudy (anonymous):

acres

OpenStudy (pfenn1):

k must be a numeral (but you are correct that it is the number of acres per year). Just look at these equations numerically. Numerically what must k equal?

OpenStudy (anonymous):

ok, you are doing everything right to help me and i am just not clicking it.... grrr... k=45?

OpenStudy (anonymous):

f(x)=15-45 is that close to the answer?

OpenStudy (pfenn1):

The equation that we are trying to solve is in the form of \[f(x)=kx\]We know when x=1, f(1)=15 million=1k x=2, f(2)=30 million=2k etc. or we could rearrange the original equation f(x)=kx or\[k=\frac{f(x)}x\]In this case, we can rearrange all the equations above so that k=f(1)/1= 15 million/1 k=f(2)/2 = 30 million/2 etc. So what must k be?

OpenStudy (anonymous):

135

OpenStudy (pfenn1):

k = 15 million/1 = 30 million/2 = 45 million/3...... therefor k = 15 million. f(x)=15,000,000x

OpenStudy (pfenn1):

Does that make sense?

OpenStudy (anonymous):

wow, I was way off ....this math is gonna be the death of me. I get SOME nd then not some. THANK YOU f or your patience. :)

OpenStudy (pfenn1):

No problem. I'm impressed that you are studying all this if you are really the mom of a teenager and a toddler. You have your hands full.

OpenStudy (anonymous):

I have a 3 year old and a 18 yr old, I am in school full-time trying for my business degree to enhance my career for myself and my family as I work full time as well. So I do certainly have my hands full LOL ... THANK YOU AGAIN. :)

OpenStudy (pfenn1):

I have 21 year old and 18 (next month) year old and I graduated college 34 years ago so sometimes I don't remember how to do some of these problems. But I try to help. Good luck with your schooling.